In acceleration of a body is a=20t^3 -3t^2+6. Find the displacement after 2.7s assuming the initial velocity is zero and the initial displacement is 4m
Answer #1

= ( Givena acceleration of body (a) = 2073_37²+6 m/s2 initial velocity = 0 = att zo ,u=0 mis initial displacement = 4m =) at= dne where MA - velocity can be writtenced, o= du wheren=displacement is done = 544-43 +6+ of du= (5t-t3 +67). dt Integrati